OpenStudy (anonymous):

cos(2arcsin x) write as an algebraic expression of x

OpenStudy (anonymous):

did you see my answer to the last one?

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

this is a completely different question though

OpenStudy (anonymous):

for this one you need \[\cos(2\theta)=1-2\sin^2(x)\]

OpenStudy (anonymous):

oops make that \[\cos(2\theta)=1-2\sin^2(\theta)\]

OpenStudy (anonymous):

replace \(\theta\) by \(\arcsin(x)\) and you are done in one step pretty much

OpenStudy (anonymous):

let me know if that is not clear, i can write it out

OpenStudy (anonymous):

alright ill see how i do with that. thanks ^.^

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

yes, that is the first step but what is \[\sin(\arcsin(x))\]?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

but sin is squared?

OpenStudy (anonymous):

ok then it should be clear if you understand \[\sin^2(\arcsin(x))\] means \[(\sin(\arcsin(x))^2\]

OpenStudy (anonymous):

1-2x^2

OpenStudy (anonymous):

youre awesome!!! thank you soo much!!! <3

OpenStudy (anonymous):

aahhh i thought that might be the confusion \[\sin^2(\theta)\] means \((\sin(\theta))^2\) yes you have it

OpenStudy (anonymous):

yupp i got it right!

OpenStudy (anonymous):

yay yw

OpenStudy (anonymous):

thanks to your help! goodnight

OpenStudy (anonymous):

gnight